OpenStudy (anonymous):

i need help who to solve √45 + √5

OpenStudy (anonymous):

can you help me

OpenStudy (anonymous):

plzzz

OpenStudy (anonymous):

yes hold on

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

woulndt you add the sqaure root of each on a calculater

OpenStudy (anonymous):

idk how to solve

OpenStudy (anonymous):

Ooh all you gotta do is what are is a factor of 45 that can be sqrt? 9 and 9 * 5 = 45 So you sqrt 9 = 3 So your answer = SQRT 3 SQRT 5 I think.. havent done it in a while

OpenStudy (anonymous):

addition

OpenStudy (anonymous):

the answer is \[4\sqrt{5}\]

OpenStudy (anonymous):

oh wait and then you have to do the other 5.... idk anymore ^.^ That work was only for the SQRT 45

OpenStudy (anonymous):

that is the answer

OpenStudy (anonymous):

good job :D

OpenStudy (anonymous):

\[\surd 45+\surd 5=3 \sqrt{5}+\surd 5=4 \sqrt{5}\]

OpenStudy (anonymous):

thanks very much